ELECTRONIC SPECTRA OF A DISCHARGE PRODUCT OF CARBON DISULFIDE: DETECTION OF A NEW IONIC SPECIES

\underline{MASAKAZU NAKAJIMA}, YU YONEDA, YOSHIHIRO SUMIYOSHI, and YASUKI ENDO , Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Tokyo 153-8902, Japan.

New electronic transition has been observed in the 18400−19800 cm^{−1} region by discharging the CS_{2} molecule in a supersonic jet using a laser induced fluorescence technique. Since the vibronic bands completely disappear when the C_{2}H_{2} sample is used, the spectral carrier contains the sulfur atom. Rotationally resolved spectra were obtained by highresolution laser scans. The band type of all the bands is assigned to ^{2}Π_{3/2} −^{2} Π_{3/2}, and the effective rotational constant of the ground electronic state has been determined to be 0.05047(4) cm^{−1}. We concluded that the ground state of the radical is ^{2}Π_{i}, and the spectral carrier may be an ionic species because both the carbon and sulfur atoms have even numbers of electrons. Based on the results of ab initio calculations, it is considered that the negative ion, SCCS^{−}, is the most probable candidate for the spectral carrier.